# Inference for Regression with

### 23 Feb 2024 v1

@@ 1 Introduction
As the amount of digitally recorded unstructured data continues to grow rapidly, empirical work in economics is increasingly incorporating it. The leading example of such data is text (Gentzkow et al., 2019a , Ash and Hansen, 2023), but others include surveys, images, and audio recordings. One of the primary applications of such data is to recover some latent variable of economic interest with an information retrieval (IR) model. Examples abound: Baker et al., (2016) measures economic policy uncertainty with newspaper text; Hoberg and Phillips, (2016) infers firms’ latent industries with corporate filings; Hansen et al., (2018) constructs measures of policy deliberation from Federal Open Market Committee (FOMC) transcripts; Magnolfi et al., (2022) uses survey data to measure product differentiation; Compiani et al., (2023) measures substitutability between products using Amazon text and image data; Gorodnichenko et al., (2023) measures tone-of-voice from audio recordings of FOMC press conferences; Gabaix et al., (2023) imputes firm characteristics from investor holdings data; Einav et al., (2022) infer patients’ health status from survey answers; Vafa et al., (2023) constructs a measure of labor market experience based on CVs. These derived measures are rarely an end in themselves. Rather, the motivation for constructing them is to study how the concept they proxy interacts with the economic environment. As such, they are typically plugged into downstream econometric models whose parameters are the main object of study. Importantly, the IR and econometric models are almost always treated as wholly separate, with the output of the former treated as “data” in the latter. We call this the two-step strategy.
@@ 2.2 Two-Step Strategy
because $\mathbb{E}[\hat{\theta}_{i}|\theta_{i},C_{i}]=\theta_{i}$ and $\mathrm{Var}(\hat{\theta}_{i})=\mathrm{Var}(\theta_{i})+\mathbb{E}[C_{i}^{-1}]\mathbb{E}[\theta_{i}(1-\theta_{i})]$ by the law of total variance and independence of $C_{i}$ and $\theta_{i}$. Evidently, there is an attenuation bias caused by measurement error in $\hat{\theta}_{i}$ which makes $\hat{\gamma}_{1}$ inconsistent.
@@ 3 Full Analysis of the Two-Step Strategy
One limiting feature of the stylized model in the previous section is that the observed data are not high-dimensional. In this section, we relax this and allow each unstructured data observation to lie in a high-dimensional space which requires some dimensionality reduction prior to regression analysis. In this section, we first describe the statistical framework linking unstructured data and the downstream regression model. We then analyze the usual two-step strategy and shed light on when it leads to valid inference, extending the findings from the stylized model in Section 2 to a general setting. In particular, two-step inference is valid only when the amount of unstructured data per observation is much larger than sample size, so that measurement error is of smaller order than sampling error. Otherwise, confidence intervals based on the usual two-step strategy have the correct width but incorrect centering, and therefore have coverage rates below nominal coverage. In the next section, we supplement the statistical framework with an upstream model linking covariates to the unstructured data to produce the Supervised Topic Model with Covariates and discuss how it solves some of these inference problems.
@@ 3.1 Statistical Framework
We begin by specifying a statistical model that, broadly speaking, has two parts. The first part computes low-dimensional numerical representations of the unstructured data. The second part introduces these numerical representations as covariates, potentially along with other quantitative data, into a linear regression model. There are a wide array of methods for dimensionality reduction used in the literature. We focus on factor modeling, which, in the context of high-dimensional discrete data, is also called topic modeling. We make this choice for two reasons. First, topic models have a well-defined statistical structure which facilitates theoretical analysis. Second, there is a large empirical literature which uses topic models as part of the two-step strategy, most commonly using textual data. Examples include Hansen et al., (2018), Mueller and Rauh, (2018), Larsen and Thorsrud, (2019), Thorsrud, (2020), Bybee et al., (2020), and Adams et al., (2021). However, the use of topic models is not limited to textual data. For example, Bandiera et al., (2020), Draca and Schwarz, (2021), and Munro and Ng, (2022) use topic models to analyze survey data. Meanwhile, Nimczik, (2017) and Olivella et al., (2021) use topic models for network-structured data.
@@ 3.1.1 Model
We consider a setting where each unstructured observation $i$ is described by $\mathbf{x}_{i}$, a $V$-dimensional vector of count variables, where $x_{i,v}$ is the number of times a feature $v$ appears in observation $i$. We consider $V$ to be high dimensional. This setting is not overly restrictive, as many types of unstructured data are naturally high-dimensional and discrete. For example, in the bag-of-words model $V$ is the number of unique terms in a textual corpus, typically in the thousands, and $x_{i,v}$ is the count of term $v$ in document $i$. The first part of the model generates a $K$-dimensional representation $\bm{\theta}_{i}$ of $\mathbf{x}_{i}$, where $K\ll V$. The second part introduces these low-dimensional representations as covariates, potentially along with other quantitative data $\mathbf{q}_{i}$, into a linear regression model:
$Y_{i}=\bm{\gamma}^{T}\bm{\theta}_{i}+\bm{\alpha}^{T}\mathbf{q}_{i}+\varepsilon_{i},\quad\quad\mathbb{E}\left[\left.\varepsilon_{i}\right|\bm{\theta}_{i},\mathbf{q}_{i}\right]=0.$ (4)
In most empirical applications in economics and finance, the key parameters of interest are the $\bm{\gamma}$ regression coefficients. Hence, we will focus mainly on estimation and inference for $\bm{\gamma}$ in what follows.
The model we consider for the unstructured data is widely used in practice but also tractable enough that we can develop theory for why the two-step strategy leads to biased inference in empirically realistic settings. As $\mathbf{x}_{i}$ is a vector of counts, it is without loss of generality to model it as a Multinomial distribution. We impose additional structure on the count probabilities for interpretability. The model is based on Probabilistic Latent Semantic Analysis (Hofmann, 1999, PLSA), a widely used factor model for discrete data, and its close cousin Latent Dirichlet Allocation (Blei et al., 2003, LDA). Formally,
@@
where $C_{i}=\sum_{v=1}^{V}x_{i,v}$ is the count of all features in observation $i$—a measure of the amount of unstructured data for observation $i$—and the count probabilities have a factor structure $\mathbf{p}_{i}=\mathbf{B}^{T}\bm{\theta}_{i}$. [^10] This model nests as a special case a pure multinomial model where $\mathbf{B}=\mathbf{I}$ and $\bm{\theta}_{i}=\mathbf{p}_{i}$. There are $K$ separate distributions over the $V$ features denoted $\bm{\beta}_{1},\ldots,\bm{\beta}_{K}$ where each $\bm{\beta}_{k}$ lies in the $(V-1)$-dimensional simplex. In text applications, these distributions are called topics, but more generally they represent common factors from which individual observations are built. We collect the factors into a $K\times V$ row-stochastic matrix $\mathbf{B}$ where $\mathbf{B}^{T}=[\bm{\beta}_{1},\ldots,\bm{\beta}_{K}]$. Each observation $i$ is characterized by the latent vector $\bm{\theta}_{i}$ which lies in the $(K-1)$-dimensional simplex. Its elements $\theta_{i,k}$ represent the weight attached to $\bm{\beta}_{k}$ in generating $\mathbf{x}_{i}$. Hence, the count probabilities for observation $i$ are $\mathbf{p}_{i}=\sum_{k=1}^{K}\mathbf{\beta}_{k}\theta_{i,k}$. It is helpful to think of $\mathbf{B}$ as a matrix of common parameters and $\bm{\theta}_{i}$ as an observation-specific latent random vector. The quantity $C_{i}$ determines the degree of precision with which we can infer $\bm{\theta}_{i}$ from $\mathbf{x}_{i}$. The interplay between the distribution of $C_{i}$ and the number of observations $n$ plays an important role in our theory below.
@@
Suppose each unstructured observation is a monetary policy speech. One distribution $\bm{\beta}_{k}$ might put high weight on words like ‘inflation’, ‘prices’, and ‘cpi’, so $\bm{\beta}_{k}$ would have an interpretation as price rises. The corresponding $\theta_{i,k}$ then represents how much speech $i$ discuss price rises. One research question might ask how attention paid to price rises, along with other economic conditions captured by other topics, affects policy actions. This could be captured by the $\bm{\gamma}$ coefficients in (4) where $Y_{i}$ is the policy action of speaker $i$ and $\mathbf{q}_{i}$ measures quantitative information like market forecasts for growth and inflation at the time the speech was made.
@@ 3.1.2 Data and Maintained Assumptions
The data available are a random sample $(Y_{i},\mathbf{q}_{i},\mathbf{x}_{i},C_{i})_{i=1}^{n}$ satisfying (4) and (5). We further assume that $C_{i}$ is independent of $\bm{\theta}_{i}$, $\mathbf{q}_{i}$, and $Y_{i}$, and that $Y_{i}$ and $\mathbf{x}_{i}$ are independent conditional on $C_{i}$ and $\bm{\theta}_{i}$. We emphasize that these restrictions are not essential and are made to simplify the following derivation. Our theory can easily be extended to time-series models, such as where (5) is replaced by a vector autoregression. We do not do so here, however, as the main take-aways are most clearly illustrated in the IID case.
We also assume that $\mathbf{B}$ and the $\bm{\theta}_{i}$ are identified in the sense that there is a unique decomposition $\mathbf{P}=\mathbf{B}^{T}\bm{\Theta}$ with $\mathbf{P}=[\mathbf{p}_{1},\ldots,\mathbf{p}_{n}]$ collecting the vectors of count probabilities across observations and $\bm{\Theta}=[\bm{\theta}_{1},\ldots,\bm{\theta}_{n}]$ collecting the topic weights across observations. For instance, identification is commonly achieved in text applications by assuming the existence of certain anchor words: these are words that are known to appear in certain topics but not others. We impose this identifiability condition because our objective is to analyze the consequences of the two-step inference approach in a transparent way. Adding partial identification into the mix will significantly complicate the analysis but may be an interesting extentsion in future research.
@@ 3.2 Theory for the Two-Step Strategy
    $Y_{i}$ is regressed on $\hat{\bm{\theta}}_{i}$ and $\mathbf{q}_{i}$. Conventional OLS standard errors are reported, treating the $\hat{\bm{\theta}}_{i}$ as if they are regular numeric data.
Evidently there is a measurement error problem: the estimates $\hat{\bm{\theta}}_{i}$ are noisy proxies for the true $\bm{\theta}_{i}$ appearing in the regression model (4). But Step (ii) overlooks this problem and treats the first-stage estimates $\hat{\bm{\theta}}_{i}$ as regular numeric data. This raises the possibility that OLS estimates of $\bm{\gamma}$ may be biased due to measurement error introduced in Step (i). Moreover, conventional standard errors are typically reported for inference on $\bm{\gamma}$. These do not account for any additional variation introduced by using noisy $\hat{\bm{\theta}}_{i}$ instead of $\bm{\theta}_{i}$, raising the possibility that inference may be biased.
In this section, we use the above statistical framework to explore when this approach delivers valid estimates and inference for $\bm{\gamma}$. To focus on the key conceptual issues, we abstract away from any additional covariates $\mathbf{q}_{i}$ in the regression equation (4). [^11] This is not restrictive, as any additional numeric covariates can be partialled-out at the cost of more complicated notation. Similarly, the following analysis and findings extend easily to models where (4) is replaced by $Y_{i}=\bm{\gamma}^{T}(\mathbf{A}\bm{\theta}_{i})+\bm{\alpha}^{T}\mathbf{q}_{i}+\varepsilon_{i}$ for some known matrix $\mathbf{A}$. See Section 3.3. Once $\mathbf{q}_{i}$ is omitted the regression still contains an intercept because the elements of $\bm{\theta}_{i}$ sum to one. In this simplified setting, the OLS estimator of $\bm{\gamma}$ is given by
$\hat{\bm{\gamma}}=\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}Y_{i}\right).$ (6)
@@ 3.2.1 Fixed Population
We first consider the large-sample properties of $\hat{\bm{\gamma}}$ where the number of observations becomes large ($n\to\infty$) but the distribution of $(Y_{i},\mathbf{q}_{i},\mathbf{x}_{i},C_{i})_{i=1}^{n}$ is held fixed. This fixed-population asymptotic framework captures a setting where the amount of unstructured data per observation is small relative to the overall sample size, as commonly encountered in empirical work.
There are many different ways of estimating $\mathbf{B}$ and $\bm{\Theta}:=[\bm{\theta}_{1},\ldots\bm{\theta}_{n}]$ in (5). For instance, one could use LDA (Blei et al., 2003) or more recent methods developed by Bing et al., (2020), Wu et al., (2023), Ke and Wang, (2022), and many others. As our objective is to focus on the consequences of the above two-step strategy, we abstract from algorithmic-specific details and instead impose some mild high-level conditions on the estimators $\hat{\mathbf{B}}$ of $\mathbf{B}$ and $\hat{\bm{\Theta}}$ of $\bm{\Theta}$. Let $\hat{\mathbf{P}}=[\hat{\mathbf{p}}_{1},\ldots,\hat{\mathbf{p}}_{n}]=[\mathbf{x}_{1}/C_{1},\ldots,\mathbf{x}_{n}/C_{n}]$ denote the $V\times n$ matrix of sample frequencies (or term frequencies, in text applications). Let $\to_{p}$ denote convergence in probability as the number of observations $n$ becomes large.
@@
    $\mathbf{B}$ and $\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]$ have full rank.
@@
Assumption 1(i) says that there are no fewer than $K$ topics. We view this as a weak restriction as $K$ is typically much smaller than $V$ in applications. Assumption 1(ii) says that the estimator $\hat{\mathbf{B}}$ is consistent for the topic weights $\mathbf{B}$. This is a mild condition satisfied by many estimators for topic models. Assumption 1(iii) imposes some structure on the estimators $\hat{\bm{\theta}}_{i}$ that we leverage to derive the asymptotic properties of $\hat{\bm{\gamma}}$. Note that Assumption 1(iii) is not vacuous: we have $\bm{\theta}_{i}=(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\mathbf{p}_{i}$ (by Assumption 1(i)) so, given any consistent estimator $\hat{\mathbf{B}}$ of $\mathbf{B}$, one could estimate $\bm{\theta}_{i}$ simply by setting $\hat{\bm{\theta}}_{i}=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}$. In that case, $\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|=0$. Also note that the first condition in (i) and parts (ii) and (iii) hold trivially for the pure multinomial model because $\hat{\mathbf{B}}=\mathbf{B}=\mathbf{I}$ and $\hat{\mathbf{p}}_{i}=\hat{\bm{\theta}}_{i}$.
Our first main result shows that the OLS estimator of $\bm{\gamma}$ in equation (6) is inconsistent in this fixed-population setting. Let $\mathrm{diag}(\mathbf{v})$ denote a diagonal matrix whose diagonal elements are the elements of the vector $\mathbf{v}$.

### 24 Mar 2024 v2

@@ 1 Introduction
As the amount of digitally recorded unstructured data continues to grow rapidly, empirical work in economics is increasingly incorporating it. The leading example of such data is text (Gentzkow et al., 2019a , Ash and Hansen, 2023), but others include surveys, images, and audio recordings. One of the primary applications of such data is to recover some latent variable of economic interest with an information retrieval (IR) model. Examples abound: Baker et al., (2016) measures economic policy uncertainty with newspaper text; Hoberg and Phillips, (2016) infers firms’ latent industries with corporate filings; Hansen et al., (2018) constructs measures of policy deliberation from Federal Open Market Committee (FOMC) transcripts; Magnolfi et al., (2022) uses survey data to measure product differentiation; Compiani et al., (2023) measures substitutability between products using Amazon text and image data; Gorodnichenko et al., (2023) measures tone-of-voice from audio recordings of FOMC press conferences; Gabaix et al., (2023) imputes firm characteristics from investor holdings data; Einav et al., (2022) infers patients’ health status from survey answers; Vafa et al., (2023) constructs a measure of labor market experience based on CVs. These derived measures are rarely an end in themselves. Rather, the motivation for constructing them is to study how the concept they proxy interacts with the economic environment. As such, they are typically plugged into downstream econometric models whose parameters are the main object of study. Importantly, the IR and econometric models are almost always treated as wholly separate, with the output of the former treated as “data” in the latter. We call this the two-step strategy.
@@ 2.2 Two-Step Strategy
because $\mathbb{E}[\hat{\theta}_{i}|\theta_{i},C_{i}]=\theta_{i}$ and $\mathrm{Var}(\hat{\theta}_{i})=\mathrm{Var}(\theta_{i})+\mathbb{E}[C_{i}^{-1}]\mathbb{E}[\theta_{i}(1-\theta_{i})]$ by the law of total variance and independence of $C_{i}$ and $\theta_{i}$. Evidently, there is an attenuation bias caused by measurement error in $\hat{\theta}_{i}$ which makes $\hat{\gamma}_{1}$ inconsistent. .
@@ 3 Full Analysis of the Two-Step Strategy
One limiting feature of the stylized model in the previous section is that the observed data are not high-dimensional. In this section, we allow each unstructured observation to lie in a high-dimensional space, which requires some dimensionality reduction prior to regression analysis. We first describe the statistical framework linking unstructured data and the downstream regression model. We then analyze the two-step strategy and show why it leads to biased inference in empirically plausible settings. In the next section, we discuss how the one-step strategy solves these inference problems.
@@ 3.1 Statistical Framework
We begin by specifying a statistical model that, broadly speaking, has two parts. The first part computes low-dimensional numerical representations of the unstructured data. The second part introduces these numerical representations as covariates, potentially along with other quantitative data, into a linear regression model. There are a wide array of methods for dimensionality reduction used in the literature. We focus on factor modeling, which, in the context of high-dimensional discrete data, is also called topic modeling. We make this choice for two reasons. First, topic models have a well-defined statistical structure which facilitates theoretical analysis. Second, there is a large empirical literature which uses topic models as part of the two-step strategy, most commonly using textual data. Examples include Hansen et al., (2018), Mueller and Rauh, (2018), Larsen and Thorsrud, (2019), Thorsrud, (2020), Bybee et al., (2020), Adams et al., (2021), and Ash et al., (2022). However, the use of topic models is not limited to textual data. For example, Bandiera et al., (2020), Draca and Schwarz, (2021), and Munro and Ng, (2022) use topic models to analyze survey data. Meanwhile, Nimczik, (2017) and Olivella et al., (2021) use topic models for network-structured data.
@@ 3.1.1 Model
Many types of unstructured data are high-dimensional and discrete. We therefore let each unstructured observation $i$ be described by $\mathbf{x}_{i}$, a $V$-dimensional vector of count variables. The value $x_{i,v}$ is the number of times a feature $v$ appears in observation $i$. We allow $V$ to be high dimensional. For example, in the bag-of-words model $V$ is the number of unique terms in a textual corpus, typically in the thousands, and $x_{i,v}$ is the count of term $v$ in document $i$.
The first part of the model generates a $K$-dimensional representation $\bm{\theta}_{i}$ of $\mathbf{x}_{i}$, where $K\ll V$. The second part introduces these low-dimensional representations as covariates, potentially along with other quantitative data $\mathbf{q}_{i}$, into a linear regression model:
$Y_{i}=\bm{\gamma}^{T}(\bm{S}\bm{\theta}_{i})+\bm{\alpha}^{T}\mathbf{q}_{i}+\varepsilon_{i},\quad\quad\mathbb{E}\left[\left.\varepsilon_{i}\right|\bm{\theta}_{i},\mathbf{q}_{i}\right]=0,$ (4)
where $\bm{S}$ is a $H\times K$ selection matrix which is chosen by the researcher. Each row of $\bm{S}$ selects a component of $\bm{\theta}_{i}$ to be included in the regression. In most applications in economics and finance, $\bm{\gamma}$ is the key parameter of interest. In other applications, however, $\bm{\alpha}$ is the focus and $\bm{S}\bm{\theta}_{i}$ plays the role of a text-derived control variable.
The model we consider for the unstructured data is widely used in practice and tractable enough that we can develop clean characterizations for the two-step strategy. As $\mathbf{x}_{i}$ is a vector of counts, it is without loss of generality to model it as Multinomial. We impose some structure on the count probabilities for interpretability. The model is based on Probabilistic Latent Semantic Analysis (Hofmann, 1999, PLSA), a widely used factor model for discrete data, and its close cousin Latent Dirichlet Allocation (Blei et al., 2003, LDA). Formally,
@@
where $C_{i}=\sum_{v=1}^{V}x_{i,v}$ is the count of all features in observation $i$—a measure of the amount of unstructured data for observation $i$—and the count probabilities $\mathbf{p}_{i}$ have the factor structure $\mathbf{p}_{i}=\mathbf{B}^{T}\bm{\theta}_{i}$. This model nests as a special case a pure multinomial model where $\mathbf{B}=\mathbf{I}$ and $\bm{\theta}_{i}=\mathbf{p}_{i}$.
There are $K$ separate distributions over the $V$ features denoted $\bm{\beta}_{1},\ldots,\bm{\beta}_{K}$ where each $\bm{\beta}_{k}$ lies in the $(V-1)$-dimensional simplex. The distributions are called topics in text applications. More generally, they represent common factors from which individual observations are built. We collect the factors into a $K\times V$ row-stochastic matrix $\mathbf{B}$ where $\mathbf{B}^{T}=[\bm{\beta}_{1},\ldots,\bm{\beta}_{K}]$. Each observation $i$ is characterized by the latent vector $\bm{\theta}_{i}$ which lies in the $(K-1)$-dimensional simplex. Its elements $\theta_{i,k}$ represent the weight attached to $\bm{\beta}_{k}$ in generating $\mathbf{x}_{i}$. Hence, the count probabilities for observation $i$ are $\mathbf{p}_{i}=\sum_{k=1}^{K}\bm{\beta}_{k}\theta_{i,k}$. Note that $\mathbf{B}$ is a matrix of common parameters whereas $\bm{\theta}_{i}$ is an observation-specific latent random vector. The quantity $C_{i}$ determines the degree of precision with which we can infer $\bm{\theta}_{i}$ from $\mathbf{x}_{i}$. The interplay between $C_{i}$ and the number of observations $n$ plays an important role in our theory.
@@
Suppose each unstructured observation is a monetary policy speech. One distribution $\bm{\beta}_{k}$ might put high weight on words like ‘inflation’, ‘prices’, and ‘cpi’, so $\bm{\beta}_{k}$ would have an interpretation as price rises. The corresponding $\theta_{i,k}$ represents how much speech $i$ discuss price rises. One research question might ask how attention paid to price rises, along with other economic conditions captured by other topics, affects policy actions. This could be captured by the $\bm{\gamma}$ coefficients in (4) where $Y_{i}$ is the policy action of speaker $i$, $\bm{S}$ selects the price rises topic weight from $\bm{\theta}_{i}$ and discards irrelevant topics (e.g., words used in generic conversation), and $\mathbf{q}_{i}$ measures quantitative information like market forecasts for growth and inflation at the time the speech was made.
@@ 3.1.2 Data and Maintained Assumptions
The data are a random sample $(Y_{i},\mathbf{q}_{i},\mathbf{x}_{i},C_{i})_{i=1}^{n}$ satisfying (4) and (5). To simplify derivations, we further assume that $C_{i}$ is independent of $(\bm{\theta}_{i},\mathbf{q}_{i},Y_{i})$, and that $(\mathbf{q}_{i},Y_{i})$ and $\mathbf{x}_{i}$ are independent conditional on $(C_{i},\bm{\theta}_{i})$. We also assume that $\mathbf{B}$ is identified. That is, there is a unique decomposition $\mathbf{P}=\mathbf{B}^{T}\bm{\Theta}$ with $\mathbf{P}=[\mathbf{p}_{1},\ldots,\mathbf{p}_{n}]$ collecting the vectors of count probabilities across observations and $\bm{\Theta}=[\bm{\theta}_{1},\ldots,\bm{\theta}_{n}]$ collecting the topic weights across observations. w Identification is commonly achieved in text applications by assuming the existence of anchor words that are known to appear in some topics but not others. We assume identifiability because our objective is to analyze the consequences of the two-step inference approach in a transparent way. Adding partial identification will significantly complicate the analysis but may be an interesting extension in future research.
@@ 3.2 Theory for the Two-Step Strategy
    $Y_{i}$ is regressed on $\bm{S}\hat{\bm{\theta}}_{i}$ and $\mathbf{q}_{i}$. Inference is performed treating the $\hat{\bm{\theta}}_{i}$ as if they are regular numeric data.
Evidently there is a measurement error problem: the estimates $\hat{\bm{\theta}}_{i}$ are noisy proxies for the true $\bm{\theta}_{i}$ appearing in the regression model (4). But Step (ii) overlooks this problem and treats the first-stage estimates $\hat{\bm{\theta}}_{i}$ as regular numeric data. This raises the possibility that OLS estimates of $\bm{\psi}$ may be biased due to measurement error introduced in Step (i). Moreover, conventional standard errors are typically reported. But these do not account for any additional variation introduced by using $\hat{\bm{\theta}}_{i}$ instead of $\bm{\theta}_{i}$.
We now analyze the two-step strategy and show how it can lead to biased estimates and inference. Let
$\bm{\psi}=\left[\begin{array}[]{c}\bm{\gamma}\\ \bm{\alpha}\end{array}\right],\quad\quad\bm{\xi}_{i}=\left[\begin{array}[]{c}\bm{S}\bm{\theta}_{i}\\ \mathbf{q}_{i}\end{array}\right],\quad\quad\hat{\bm{\xi}}_{i}=\left[\begin{array}[]{c}\bm{S}\hat{\bm{\theta}}_{i}\\ \mathbf{q}_{i}\end{array}\right].$
The OLS estimator of $\bm{\psi}$ in the two-step strategy is given by
$\hat{\bm{\psi}}=\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}Y_{i}\right).$ (6)
@@ 3.2.1 Fixed Population
We first consider the large-sample properties of $\hat{\bm{\psi}}$ where the number of observations becomes large ($n\to\infty$) but the distribution of $(Y_{i},\mathbf{q}_{i},\mathbf{x}_{i},C_{i})_{i=1}^{n}$ is held fixed. This asymptotic framework mimics many empirical designs where there is a relatively small amount of unstructured data per observation but a large number of observations.
There are many different ways of estimating $\mathbf{B}$ and $\bm{\theta}_{i}$ in (5). For instance, one could use LDA (Blei et al., 2003) or more recent methods developed by Bing et al., (2020), Wu et al., (2023), Ke and Wang, (2022), and many others. As our objective is to focus on the consequences of the two-step strategy, we abstract from algorithmic-specific details and instead impose mild conditions on the estimators $\hat{\mathbf{B}}$ of $\mathbf{B}$ and $\hat{\bm{\theta}}_{1},\ldots,\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{1},\ldots,\bm{\theta}_{n}$. Let $\hat{\mathbf{p}}_{i}=\mathbf{x}_{i}/C_{i}$, $i=1,\ldots,n$. Let $\to_{p}$ denote convergence in probability as $n\to\infty$.
@@
    $\mathbf{B}$ has full rank.
@@
  4. (iv)
    $\mathbb{E}\left[\|\mathbf{q}_{i}\|^{2}\right]<\infty$ and $\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ has full rank.
Assumption 1(i) says that there are no fewer than $K$ topics. This is a weak restriction, as $K$ is typically much smaller than $V$ in applications. Assumption 1(ii) says that $\hat{\mathbf{B}}$ is consistent for the topic weights $\mathbf{B}$. This is a mild condition satisfied by many estimators for topic models. Assumption 1(iii) imposes some structure on the estimators $\hat{\bm{\theta}}_{i}$. This condition is not vacuous: we have $\bm{\theta}_{i}=(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\mathbf{p}_{i}$ (by Assumption 1(i)) so, given any consistent estimator $\hat{\mathbf{B}}$ of $\mathbf{B}$, one could estimate $\bm{\theta}_{i}$ simply by setting $\hat{\bm{\theta}}_{i}=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}$. In that case, $\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|=0$. Assumption 1(i)-(iii) holds trivially for the pure multinomial model because $\hat{\mathbf{B}}=\mathbf{B}=\mathbf{I}$ and $\hat{\mathbf{p}}_{i}=\hat{\bm{\theta}}_{i}$. Finally, Assumption 1(iv) imposes a standard moment condition on the numeric regressors $\mathbf{q}_{i}$ and ensures that these are not perfectly collinear with the included latent variables $\bm{S}\bm{\theta}_{i}$.
Our first main result shows that $\hat{\bm{\psi}}$ is inconsistent when the amount of unstructured data per observation is small relative to the sample size. Let $\mathrm{diag}(\mathbf{v})$ denote a diagonal matrix whose diagonal elements are the elements of the vector $\mathbf{v}$. Let $\mathbf{Q}_{\mathbf{B}}=(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}$ and let $\mathbf{0}$ denote a conformable matrix of zeros.